# On the resolution of categorical symmetries in (Non-) Unitary Rational CFTs

### v1

@@ Acknowledgments
We would like to thank Brandon Rayhaun for detailed comments on the previous version of the manuscript. A.B and J.S would also like to thank Arpit Das and Thomas Quella for discussions on related topics and comments on some parts of the draft, as well as making us aware of the upcoming work Das:2025xyz. J.S is indebted to Suresh Govindarajan, Chethan N. Gowdigere, and Akhila S. for useful discussions. S.G (PMRF ID: 1702711) and S.P (PMRF ID: 1703278) is supported by the Prime Minister’s Research Fellowship of the Government of India. S.G would like to thank Institut des Hautes Études Scientifiques (IHES), France, for their kind hospitality during the course of the work. A.B thanks the participants of the BIRS-CMI workshop “Quantum Gravity and Information Theory: Modern Developments” (https://www.birs.ca/events/2025/5-day-workshops/25w5386) held at CMI, Chennai, during 9-14th November 2025 where part of the work was presented by J.S. J.S would like to thank the organizers of the BIRS-CMI workshop 2025 for giving warm hospitality. J.S is supported by the project CRG/2023/001120 and OTH/SERB/13451 (50:50 Postdoctoral scheme of IIT Gandhinagar). A.B is supported by the Core Research Grant (CRG/2023/001120) from the Anusandhan National Research Foundation, India. A.B also acknowledges the associateship program of the Indian Academy of Science, Bengaluru.

### v2

@@ Acknowledgments
We would like to thank Brandon Rayhaun for detailed comments on the previous version of the manuscript. A.B and J.S would also like to thank Arpit Das and Thomas Quella for discussions on related topics and comments on some parts of the draft, as well as making us aware of the upcoming work Das:2025xyz. J.S is indebted to Suresh Govindarajan, Chethan N. Gowdigere, and Akhila S. for useful discussions. S.G would like to thank Jayashish Das for discussion on related topic. S.G (PMRF ID: 1702711) and S.P (PMRF ID: 1703278) is supported by the Prime Minister’s Research Fellowship of the Government of India. S.G would like to thank Institut des Hautes Études Scientifiques (IHES), France, for their kind hospitality during the course of the work. A.B thanks the participants of the BIRS-CMI workshop “Quantum Gravity and Information Theory: Modern Developments” (https://www.birs.ca/events/2025/5-day-workshops/25w5386) held at CMI, Chennai, during 9-14th November 2025 where part of the work was presented by J.S. J.S would like to thank the organizers of the BIRS-CMI workshop 2025 for giving warm hospitality. J.S is supported by the project CRG/2023/001120 and OTH/SERB/13451 (50:50 Postdoctoral scheme of IIT Gandhinagar). A.B is supported by the Core Research Grant (CRG/2023/001120) from the Anusandhan National Research Foundation, India. A.B also acknowledges the associateship program of the Indian Academy of Science, Bengaluru.